# Data-adaptive structural change-point detection via isolation

### v2

@@ 1 Introduction
where $X_{t}$ are the observed data and $f_{t}$ is a one-dimensional deterministic signal with structural changes at certain points. We highlight, however, that the case when $X_{t}$ is multivariate is also discussed in our paper. We focus on the case of piecewise-constant and continuous piecewise-linear signals $f_{t}$, meaning that the changes are in the mean or the slope. However, the proposed algorithm can be extended to other, possibly more complicated, signal structures. Although detecting changes in the mean is a simple case, as noted by [7], more complex change-point problems that allow changes in properties other than the mean, can be reduced to the segmentation problem we are studying. This can be done by applying a suitable transformation to the data that reveals the changes as those in the mean of the transformed data. For example, [14] and [1] provide methods for the detection of multiple change-points in the second-order (i.e. autocovariance and cross-covariance) structure of possibly high dimensional time series, using Haar wavelets as building blocks. The signal is transformed into a piecewise-constant one with change-point locations identical to those in the original signal. We want to emphasise that the novelty of our paper does not come from the structure of changes that the method can be applied to. Upon correct choice of a contrast function, we can cover more scenarios such as the case of piecewise-quadratic signals (see [6] for more details). The importance and novelty of our work lies mainly in the data-adaptive nature of the proposed algorithm itself, which, at each step, starts searching for change-points in the areas that most prominently include one.
@@
Starting with the optimization-based methods for detecting changes in the mean when $f_{t}$ is piecewise-constant, one of the most common penalty functions is the Schwarz Information Criterion ([44]), which was used by [51] for change-point detection, assuming Gaussian random variables $\epsilon_{t}$ in (1). Under the same assumption, [52] studies an estimator based on least squares. Relaxing this assumption and allowing instead for more general exponential family distributions, [27] introduces a dynamic programming algorithm which uses maximum likelihood estimates of the location of the change-points. [39] introduces an AIC-type criterion for change-point models and [19] shows that while an AIC-like information criterion does not give a strongly consistent selection of the optimal number of change-points, a BIC-like information criterion does. [28] employs dynamic programming to guarantee that the exact global optimum, in terms of creating the segments, is found, while [30] proposes an algorithm, called PELT, whose computational cost is linear in the number of observations. [42] introduces the pDPA algorithm, which includes a pruning step towards complexity reduction. [36] combines the ideas from PELT and pDPA leading to two new algorithms, FPOP and SNIP, which have low computational complexity.
@@
In this project, we propose a data-adaptive change-point detection method, called Data Adaptive ISolation (labelled DAIS), that attempts to isolate the change-points before detection, by also taking into account the potential true locations of the change-points. This means that the algorithm does not start from a random interval, or the beginning/end of the signal, as most algorithms in the literature, but instead starts checking around a point that there is reason to believe it could be around a true change-point. The idea behind the data-adaptive nature of the algorithm, and the belief that the algorithm will start close to a change-point, arise from the fact that, for example, in the case that the unobserved true signal $f_{t}$ in (1) is piecewise-constant, taking pairwise differences between consecutive time steps will reveal a constant signal with the value 0, with spikes where the change-points occur. This means that the largest spike, in absolute value, occurs at the location of the change-point with the largest jump in the sequence, therefore the most prominent change-point. Similarly, in the case that $f_{t}$ is continuous and piecewise-linear, differencing the signal twice will again reveal a signal with spikes near to the true locations of the change-points. We try to use this fact in the observed signal, $X_{t}$. In the algorithm, we identify the location of this spike, which we will be referring to as the ‘largest difference’ for the rest of the paper, and test around it for possible change-points. A discussion about the connection between the isolation aspect of the algorithm and the location of the largest difference can be found in Section 2.3, while Section 3.3 complements this discussion with theoretical results in the case that $\epsilon_{t}$ is Gaussian. We want to emphasise that since the largest difference is only used as the starting point of the search for change-points in the data sequence, the true change-points are detected even when the largest difference is far from them. The novelty of our work lies in the data-adaptive nature of the algorithm as just described.
[…]
@@
The data-adaptive nature of DAIS is what differentiates it from its competitors. A key methodological difference with other algorithms is the way the intervals are being checked. WBS and NOT check to the left and right of the interval in which detection has already occurred, using pre-drawn intervals, while WBS2 draws new intervals. Essentially, for both algorithms, the data sequence is split into three parts, one being the interval including the change-point that has just been detected, one to its left and one to its right. After detection, ID re-starts on the left or the right of the interval where detection occurred, splitting the data sequence in two parts. DAIS, similarly to WBS and WBS2, checks intervals both to the left and right of the interval in which detection occurred. However, every time it re-starts, expanding intervals around the location of the largest difference, in absolute value, are created and thus detection of the change-point in an interval in which it is isolated occurs with high probability. The data-adaptivity also gives an advantage in terms of computational complexity, as is explained in Section 4.1.
@@
The paper is organized as follows. Section 2 provides details on the methodology of DAIS. We present a simple example followed by a detailed explanation of the algorithm and, in Section 2.3, we discuss the connection between the location of the largest difference calculated at each step of the algorithm and the guarantee that the change-point can be detected in an interval in which it is isolated. In Section 3, we provide theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations as well as some theoretical results concerning the discussion of Section 2.3. Section 4 includes an explanation of how some parameters of DAIS are selected and a note on computational complexity, which is also compared to that of competitors. In Section 5, we provide results on simulated data and a comparison to state-of-the-art competitors. Section 6 discusses extensions to the algorithm in more complicated signal structures, which include temporal correlation, relaxing the Gaussianity assumption on the noise term $\epsilon_{t}$, and multivariate data sequences. In Section 7, two examples using real world data are presented. The first one is about crime data reported daily and the second one involves weekly data on the Euro to British pound exchange rate. Section 8 concludes the paper with a summary of the most important findings. In Appendix A, we provide the signals used in the simulation study of Section 5 and in Appendix B further simulation results can be found. Finally, the proofs of Theorems 1 and 2 and the outline of the proof of Theorem 3 can be found in Appendix C, D and E, respectively.
@@ 2.1 Simple example
We emphasise that the differences are only used as a way to start the algorithm in an area close to the change-point. In cases with high variance, $\sigma^{2}$, of the noise, the largest difference might not be around the change-point but under the assumptions specified in Section 3, the change-points will still be detected with high probability. The location of the largest difference that guarantees detection of the change-point in an interval in which it is isolated is discussed in Section 2.3, and further results are provided in Section 3.3 for the case that $\epsilon_{t}$ in (1) is Gaussian.
@@ 2.2 The DAIS algorithm
$d_{s,e}=\left\{\begin{array}[]{ll}\textrm{argmax}_{t\in\{s,s+1,\dots,e-1\}}\{\lvert X_{t+1}-X_{t}\rvert\},&f_{t}\textrm{ piecewise-constant},\\ \textrm{argmax}_{t\in\{s,s+1,\dots,e-2\}}\{\lvert X_{t+2}-2X_{t+1}+X_{t}\rvert\},&f_{t}\textrm{ piecewise-linear.}\end{array}\right.$ (2)
@@
These are collected in the ordered sets, $L_{d_{s,e},s,e}$ and $R_{d_{s,e},s,e}$, defined in (2.2), which consist of the left and right end-points, respectively, of the intervals used in DAIS. Since the expansions occur in turn once from the right and once from the left, the points appear twice in each set, up until the moment when one end-point is equal to either $s$ or $e$. The left- and right-expanding intervals checked are collected in the ordered set $I_{j}$ in (4). The ordered sets of the end-points for the expansions are:
@@
Algorithm 1 provides a pseudocode that briefly explains DAIS’ workflow. The first step of DAIS is to identify, within the observed data, $X_{t}$, $t=1,2,\ldots,T$, the location of the largest difference $d_{1,T}$, as defined in (2). This is the step that makes the algorithm data-adaptive, differentiating it from the competitors, as the algorithm will check different intervals $I_{j}$ for the same underlying signals (in terms of the type of the change, the length of the sequence and the change-point locations) if they have different observed noise, depending on the value of $d_{1,T}$. Applying the appropriate transformation (as in (2) based on the signal structure) to the unobserved true signal $f_{t}$ in (1), will reveal a constant signal at 0, with non-zero values at the locations where the change-points occur, so we expect that the largest difference for the observed data $X_{t}$ will be at a point near the change-point with the largest magnitude of change. A proper explanation of what we consider to be ‘near’ is discussed in Section 2.3 in detail. After calculating $d_{1,T}$, expanding intervals are used to check for possible change-points. The algorithm expands the intervals once from the right and once from the left of the point where the maximum absolute difference occurs at every step by an expansion parameter of magnitude $\lambda_{T}$. By doing this, we identify the location of the change-point that is closest to the detected largest difference. This expansion is performed for at most $\lceil\frac{T}{\lambda_{T}}\rceil+1$ steps in total. Upon the detection of a change-point, the algorithm restarts after discarding the data that have already been checked. If, for example, the change-point was detected in the subinterval $[s^{\ast},e^{\ast}]$, then DAIS will be reapplied to $[1,s^{\ast}]$ and $[e^{\ast},T]$. When checking any interval $[s,e]$ the maximum value attained by the contrast function at a point $b$ within the interval, is compared to the chosen threshold $\zeta_{T}$. If this value exceeds $\zeta_{T}$, we take $b$ to be a change-point, for $b\in\{s,s+1,\ldots,e-1\}$ in the case of piecewise-constant signals, and $b\in\{s,s+1,\ldots,e-2\}$ in the case of piecewise-linear signals. If no change-point is detected, then the algorithm terminates.
@@
As mentioned in the introduction, the algorithm can be applied to more general signals, for example piecewise-polynomial signals of order $p$. In this case, the location of the largest difference needs to be generalised. In the more general case, it can be defined as the value of $t$ for which the $(p+1)-$times differentiated signal is maximised.
@@
There are two main advantages to the algorithm due to its data-adaptive nature. Firstly, since at each step DAIS starts checking for potential change-points around the largest difference, which as already explained we have reason to believe that it is around the most prominent change-point, then this change-point is detected fast. If the starting point is not near the change-point, detection still occurs with a slower speed. Secondly, when the change-point is at the location of the largest difference, or close to it, due to the way the expansions occur, the change-point lies in the middle of the intervals being checked when an even number of expansions have been performed. This provides advantages in detection power due to the change-point being detected in some cases in balanced intervals. Such near balance in the distances of the change-points from the left and right end-points of the intervals under consideration does not necessarily appear in many state-of-the-art competitors, such as Binary Segmentation, ID, NOT, WBS, WBS2 and SeedBS. Thus, the data-adaptive nature of the algorithm enhances DAIS’ speed and accuracy in estimating the locations of the change-points.
@@ 2.3 Location of the largest difference and the isolation aspect
As already explained in detail in Section 2.2, for a given data sequence $X_{t}$, the algorithm calculates the location of the largest difference $d_{s,e}$, as defined in (2), in the interval $[s,e]$ under consideration at each step and checks for change-points around it. For different data sequences, DAIS will start searching for change-points in different neighbourhoods of the data, based on where the most prominent change-point (if any) lies. This is exactly the novel data-adaptive nature of the proposed algorithm, compared to the existing algorithms in the literature that exhibit a fixed workflow, not adaptive to the data structure. In this subsection, we now explain the role of the location of the largest difference in the detection of a change-point in an interval where isolation is guaranteed.
As will be explained in the proof of Theorem 1 in Appendix C, the detection of a change-point $r_{j}$, for $j\in\{1,\ldots,N\},N\in\mathbb{N}$ will certainly occur if both end-points are at least at a distance $\delta_{T}/2n$ from the true location of the change-point, but of course it can happen at any smaller interval where the value of the contrast function exceeds the pre-specified threshold $\zeta_{T}$. The positive number $n$ satisfies $3/2\leq n\leq\delta_{T}/2\lambda_{T}$ where, for $r_{0}=0$ and $r_{N+1}=T$, $\delta_{T}=\min_{j=1,\ldots,N+1}\{r_{j}-r_{j-1}\}$ is the smallest distance between two consecutive change-points and $\lambda_{T}$, as explained previously, is the expansion parameter. In the worst-case scenario, meaning that the interval in which detection occurs is the biggest possible and therefore the most expansions have been performed, it holds that, firstly, both end-points are at least a distance $\delta_{T}/2n$ from the change-point and, secondly, at least one of the end-points will be in one of the intervals:
@@
Setting $n=\frac{\delta_{T}}{2\lambda_{T}}$ ensures that the detection happens as quickly as possible. Note that in [2] and [23] the choice $n=3/2$ was made while in [6] the value of $n=3$ was used when defining similar intervals for the location of the end-points of the interval where detection occurs.
@@
Since the length of the intervals in (5) is $\delta_{T}/2n$, the condition $2n\lambda_{T}\leq\delta_{T}$ is required, so that there is always at least one end-point that lies within $I^{L}_{j}$ and $I^{R}_{j}$. Since $\lambda_{T}\geq 1$, we implicitly require $\delta_{T}\geq 2n$. For $d_{s,e}\in[1,r_{1}]$ and $d_{s,e}\in[r_{N},T]$, any value of $d_{s,e}$ guarantees isolation, as the left and right expanding intervals will not include a second change-point before detection occurs, besides $r_{1}$ and $r_{N}$, respectively. Now considering the more complicated case, where the location of the largest difference lies between two change-points, define $\tilde{\delta}_{j}=r_{j+1}-r_{j}$ and $\delta_{s,e}^{j}=\min\{r_{j+1}-d_{s,e},d_{s,e}-r_{j}\}$ for $j=0,1,\ldots,N$. Without loss of generality, suppose that $d_{s,e}\in\left(r_{J},r_{J+1}\right)$ for some $J\in\{1,\ldots,N-1\}$. Isolation is guaranteed whenever the following holds
$0\leq\delta_{s,e}^{J}\leq\frac{\tilde{\delta}_{J}}{2}-\frac{3\delta_{T}}{4n}.$ (6)

### v3

@@ 1 Introduction
where $X_{t}$ are the observed data and $f_{t}$ is a one-dimensional deterministic signal with structural changes at certain points. We highlight, however, that the case when $X_{t}$ is multivariate is also discussed in our paper. We focus on the case of piecewise-constant and continuous, piecewise-linear signals $f_{t}$, meaning that the changes are in the mean or the slope. However, the proposed algorithm can be extended to other, possibly more complicated, signal structures. Although detecting changes in the mean is a simple case, as noted by [9], more complex change-point problems that allow changes in properties other than the mean, can be reduced to the segmentation problem we are studying. This can be done by applying a suitable transformation to the data that reveals the changes as those in the mean of the transformed data. For example, [16] and [2] provide methods for the detection of multiple change-points in the second-order (i.e. autocovariance and cross-covariance) structure of possibly high dimensional time series, using Haar wavelets as building blocks. The signal is transformed into a piecewise-constant one with change-point locations identical to those in the original signal. We want to emphasize that the novelty of our paper does not come from the structure of changes that the method can be applied to. Upon correct choice of a contrast function, we can cover more scenarios such as the case of piecewise-quadratic signals (see [7] for more details). The importance and novelty of our work lies mainly in the data-adaptive nature of the proposed algorithm itself, which, at each step, starts searching for change-points in the areas that most prominently include one.
@@
Starting with the optimization-based methods for detecting changes in the mean when $f_{t}$ is piecewise-constant, one of the most common penalty functions is the Schwarz Information Criterion ([54]), which was used by [63] for change-point detection, assuming Gaussian random variables $\epsilon_{t}$ in (1). Under the same assumption, [64] studies an estimator based on least squares. Relaxing this assumption and allowing instead for more general exponential family distributions, [31] introduces a dynamic programming algorithm which uses maximum likelihood estimates of the location of the change-points. [48] introduces an AIC-type criterion for change-point models and [20] shows that while an AIC-like information criterion does not give a strongly consistent selection of the optimal number of change-points, a BIC-like information criterion does. [32] employs dynamic programming to guarantee that the exact global optimum, in terms of creating the segments, is found, while [35] proposes an algorithm, called PELT, whose computational cost is linear in the number of observations. [52] introduces the pDPA algorithm, which includes a pruning step towards complexity reduction. [44] combines the ideas from PELT and pDPA leading to two new algorithms, FPOP and SNIP, which have low computational complexity. Finally, [24] introduces SMUCE, a new estimator for the change-point problem in exponential family regression and [58] proposes two procedures achieving optimal rate, a least-squares estimator with a new multiscale penalty and a two-step multiscale post-processing procedure.
@@
In this paper, we propose a data-adaptive change-point detection method, called Data Adaptive ISolation (labelled DAIS), that attempts to isolate the change-points before detection, by also taking into account the potential true locations of the change-points. This means that the algorithm does not start from a random interval, or the beginning/end of the signal, as most algorithms in the literature, but instead starts checking around a point that there is reason to believe it could be around a true change-point. The idea behind the data-adaptive nature of the algorithm, and the belief that the algorithm will start close to a change-point, arise from the fact that, for example, in the case that the unobserved true signal $f_{t}$ in (1) is piecewise-constant, taking pairwise differences between consecutive time steps will reveal a constant signal with the value 0, with spikes where the change-points occur. This means that the largest spike, in absolute value, occurs at the location of the change-point with the largest jump in the sequence, therefore the most prominent change-point. Similarly, in the case that $f_{t}$ is continuous and piecewise-linear, differencing the signal twice will again reveal a signal with spikes near to the true locations of the change-points. We try to use this fact in the observed signal, $X_{t}$. In the algorithm, we identify the location of this spike, which we will be referring to as the ‘largest difference’ for the rest of the paper, and test around it for possible change-points. A discussion about the connection between the isolation aspect of the algorithm and the location of the largest difference can be found in Section 2.3, while Section 3.3 complements this discussion with theoretical results in the case that $\epsilon_{t}$ is Gaussian. We want to emphasize that since the largest difference is only used as the starting point of the search for change-points in the data sequence, the true change-points are detected even when the largest difference is far from them. The novelty of our work lies in the data-adaptive nature of the algorithm as just described.
The DAIS algorithm uses left- and right-expanding intervals around the location of the largest difference (in absolute value) found, denoted by $d_{s,e}$, in the interval $[s,e]$ in order to identify the potential change-point that might have caused the spike and so must be close to $d_{s,e}$. This is done in a deterministic way around $d_{s,e}$, expanding once either only to the left or only to the right at each step, in an alternate way. Using expanding intervals, we achieve isolation of the change-points, which is desirable as the detection power of the contrast function is maximized in such cases (see Sections 3.1 and 3.2 for the choice of the contrast function). Due to the alternating sides of the expansions, the location of the largest difference, which, as explained before, we have a reason to believe to be around the location of a change-point, is at the midpoint of the interval being checked (or close to it) after an even number of steps has been performed. This increases the power of the contrast function and so gives an advantage to the detection power of the method. Using an expansion parameter $\lambda_{T}$ (more details in Section 4.2) at most $\lceil(e-s+1)/\lambda_{T}\rceil+1$ steps are necessary to check the whole length of the interval $[s,e]$ around the largest difference. As soon as a change-point is detected, DAIS restarts on two disjoint intervals, one ending at the start-point of the interval where the detection occurred and one starting from the end-point of that same interval. A more detailed explanation of the algorithm can be found in Section 2.
@@
The data-adaptive nature of DAIS is what differentiates it from its competitors. A key methodological difference with other algorithms is the way the intervals are being checked. WBS and NOT, as a first step, randomly draw a number of intervals, and derive the maximum of the contrast function values within those intervals. For those values exceeding a certain, predefined threshold, the change-point is detected for WBS at the location where the maximum, among the maximums within each interval, is attained and for NOT at the location corresponding to that maximum value (that surpasses the threshold) that belongs to the smallest interval drawn. WBS2 draws new random intervals every time a change-point is detected. ID uses one-sided expanding intervals starting from the end-points of the data sequence under consideration in order to achieve isolation before detection, and so the intervals are checked in a deterministic way. All aforementioned algorithms will follow the described method for any data sequence, drawing random intervals (WBS, WBS2, NOT) or starting from a fixed point (ID). In contrast to this, DAIS adapts its starting-point at every step to the nature of the data, as it takes into account the location and magnitude of the changes when calculating the location of the largest difference. Expanding intervals around the location of the largest, in absolute value, difference, are created and thus detection of the change-point in an interval in which it is isolated occurs with high probability. This way of expanding around $d_{s,e}$ increases the detection power, as explained in Section 2.2. The data-adaptivity also gives an advantage in terms of computational complexity, as is explained in Section 4.1.
@@
The paper is organized as follows. Section 2 provides details on the methodology of DAIS. We present a simple example followed by a detailed explanation of the algorithm and, in Section 2.3, we discuss the connection between the location of the largest difference calculated at each step of the algorithm and the guarantee that the change-point can be detected in an interval in which it is isolated. In Section 3, we provide theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations as well as some theoretical results concerning the discussion of Section 2.3. Section 4 includes an explanation of how some parameters of DAIS are selected and a note on computational complexity, which is also compared to that of competitors. In Section 5, we provide results on simulated data and a comparison to state-of-the-art competitors. Section 6 discusses extensions to the algorithm in more complicated signal structures, which include temporal correlation, relaxing the Gaussianity assumption on the noise term $\epsilon_{t}$, and multivariate data sequences. In Section 7, two examples using real world data are presented. The first one is about crime data reported daily and the second one involves weekly data on the Euro to British pound exchange rate. Section 8 concludes the paper with a summary of the most important findings. The proof of Theorem 1 and the outline of the proof of Theorem 3 can be found in Appendix B and C, respectively. Finally, further simulation results, an investigation of the impact of the choice of parameter values in the performance of the algorithm, and the proof of Theorem 2 can be found in the supplementary material.
@@ 2.1 Simple example
We emphasize that the differences are only used as a way to start the algorithm in an area close to the change-point. In cases with high variance, $\sigma^{2}$, of the noise, the largest difference might not be around the change-point but under the assumptions specified in Section 3, the change-points will still be detected with high probability. The location of the largest difference that guarantees detection of the change-point in an interval in which it is isolated is discussed in Section 2.3, and further results are provided in Section 3.3 for the case that $\epsilon_{t}$ in (1) is Gaussian.
@@ 2.2 The DAIS algorithm
$d_{s,e}=\left\{\begin{array}[]{ll}\underset{t\in\{s,s+1,\dots,e-1\}}{\textrm{argmax}}\{|X_{t+1}-X_{t}|\},&f_{t}\textrm{ is piecewise-constant},\\ \underset{t\in\{s,s+1,\dots,e-2\}}{\textrm{argmax}}\{\lvert X_{t+2}-2X_{t+1}+X_{t}\rvert\},&f_{t}\textrm{ is continuous, piecewise-linear.}\end{array}\right.$ (2)
@@
These are collected in the ordered sets, $L_{d_{s,e},s,e}$ and $R_{d_{s,e},s,e}$, defined in (2.2), which consist of the left and right end-points, respectively, of the intervals used in DAIS. It holds that for $m,k\leq K^{\min}$, $s\leq c_{m}^{l}<c_{m-1}^{l}$ and $c_{k-1}^{r}<c_{k}^{r}\leq e$, while for $m,k\in\{K^{\min}+1,\ldots,K^{\max}\}$, either $s=c_{m}^{l}=c_{m-1}^{l}$ and $c_{k-1}^{r}<c_{k}^{r}\leq e$ or $s\leq c_{m}^{l}<c_{m-1}^{l}$ and $c_{k-1}^{r}=c_{k}^{r}=e$ (so one end-point remains fixed). Since the expansions occur in turn once from the right and once from the left, the points appear twice in each set, up until the moment when one end-point is equal to either $s$ or $e$, which occurs when $K^{\min}$ expansions have occurred on both sides, and so $2K^{\min}$ expansions in total. The left- and right-expanding intervals checked are collected in the ordered set $I_{j}$ in (4). The ordered sets of the end-points for the expansions are:
@@
Algorithm 1 provides a pseudocode that briefly explains the workflow of DAIS. The first step is to identify, within the observed data, $X_{t}$, $t=1,2,\ldots,T$, the location of the largest difference $d_{1,T}$, as defined in (2). This is the step that makes the algorithm data-adaptive, differentiating it from the competitors, as the algorithm will check different intervals $I_{j}$ for the same underlying signals (in terms of the type of the change, the length of the sequence and the change-point locations) if they have different observed noise, depending on the value of $d_{1,T}$. Applying the appropriate transformation (as in (2) based on the signal structure) to the unobserved true signal $f_{t}$ in (1), will reveal a constant signal at 0, with non-zero values at the locations where the change-points occur, so we expect that the largest difference for the observed data $X_{t}$ will be at a point near the change-point with the largest magnitude of change. A proper explanation of what we consider to be ‘near’ is discussed in Section 2.3 in detail. After calculating $d_{1,T}$, expanding intervals are used to check for possible change-points. The algorithm expands the intervals once from the right and once from the left of the point where the maximum absolute difference occurs at every step by an expansion parameter of magnitude $\lambda_{T}$. By doing this, we identify the location of the change-point that is closest to the detected largest difference. This expansion is performed for at most $\lceil\frac{T}{\lambda_{T}}\rceil+1$ steps in total. Upon the detection of a change-point, the algorithm restarts after discarding the data that have already been checked. If, for example, the change-point was detected in the subinterval $[s^{\ast},e^{\ast}]$, then DAIS will be reapplied to $[1,s^{\ast}]$ and $[e^{\ast},T]$. When checking any interval $[s,e]$ the maximum value attained by the contrast function at a point $b$ within the interval, is compared to the chosen threshold $\zeta_{T}$. If this value exceeds $\zeta_{T}$, we take $b$ to be a change-point, for $b\in\{s,s+1,\ldots,e-1\}$ in the case of piecewise-constant signals, and $b\in\{s,s+1,\ldots,e-2\}$ in the case of continuous, piecewise-linear signals. If no change-point is detected, then the algorithm terminates.
@@
As mentioned in the Introduction, the algorithm can be applied to more general signals, for example continuous, piecewise-polynomial signals of order $p$. In this case, the location of the largest difference needs to be generalized. In the more general case, it can be defined as the value of $t$ for which the absolute value of the $(p+1)-$times differentiated signal is maximized. Differentiating the signal a $p+1$ amount of times eliminates a degree $p$ polynomial trend ([29]).
@@
There are two main advantages to the algorithm due to its data-adaptive nature. Firstly, since at each step DAIS starts checking for potential change-points around the largest difference, which as already explained we have reason to believe that it is around the most prominent change-point, then this change-point is detected fast. If the starting point is not near the change-point, detection still occurs with a slower speed. Secondly, when the change-point is at the location of the largest difference, or close to it, due to the way the expansions occur, the change-point lies in the middle of the intervals being checked when an even number of expansions have been performed. This provides advantages in detection power due to the change-point being detected in some cases in balanced intervals. Such near balance in the distances of the change-points from the left and right end-points of the intervals under consideration does not necessarily appear in many state-of-the-art competitors, such as Binary Segmentation, ID, NOT, WBS, WBS2 and SeedBS. Thus, the data-adaptive nature of our algorithm enhances its speed and accuracy in estimating the locations of the change-points.
@@ 2.3 Location of the largest difference and the isolation aspect
As already explained in detail in Section 2.2, for a given data sequence $X_{t}$, the algorithm calculates the location of the largest difference $d_{s,e}$, as defined in (2), in the interval $[s,e]$ under consideration at each step and checks for change-points around it. For different data sequences, DAIS will start searching for change-points in different neighborhoods of the data, based on where the most prominent change-point (if any) lies. This is exactly the novel data-adaptive nature of the proposed algorithm, compared to the existing algorithms in the literature that exhibit a fixed workflow, not adaptive to the data structure regarding the location of the change-points, if any exist. In this section, we now explain the role of the location of the largest difference in the detection of a change-point in an interval where isolation is guaranteed, a discussion which is related to Assumption (A1) of Section 3.1.
For the rest of this section, we are working under the framework of Assumptions (A2) and (A3), for piecewise-constant and continuous, piecewise-linear signals, respectively. The assumptions can be found in Sections 3.1 and 3.2 and need to be satisfied for the change-points to be detected. The detection of a change-point $r_{j}$, for $j\in\{1,\ldots,N\},N\in\mathbb{N}$ will certainly occur if both end-points are far enough from the true location of the change-point, but of course it can happen at any smaller interval where the value of the contrast function exceeds the pre-specified threshold $\zeta_{T}$. We define the smallest distance between two consecutive change-points by $\delta_{T}$, as in (14). The constant $n\geq 3/2$, also introduced in this section, acts as a scaling parameter and in the worst case scenario, meaning that the interval in which detection occurs is the largest possible, both end-points are at least at a distance $\delta_{T}/2n$ from the true change-point (as also explained in the proof of Theorem 1 in Appendix B). Larger values of $n$ imply that the end-points of the intervals will be closer to the change-points when detection occurs. In the worst-case scenario, as just explained, it holds that not only both end-points are at least a distance $\delta_{T}/2n$ from the change-point, but also at least one of the end-points will be in one of the intervals:
@@
Note that in [26] and [3] the choice $n=3/2$ was made while in [7] the value of $n=3$ was used when defining similar intervals for the location of the end-points of the interval which allow for detection to occur. Since the length of the intervals in (5) is $\delta_{T}/2n$, the condition $\lambda_{T}\leq\delta_{T}/2n$ is required, so that there is always at least one end-point that lies within $I^{L}_{j}$ and $I^{R}_{j}$, where, as previously explained, $\lambda_{T}$ is the expansion parameter. Since $\lambda_{T}\geq 1$, we implicitly require $\delta_{T}\geq 2n$ and, using the lower bound of $n$, $\lambda_{T}\leq\delta_{T}/3$.
@@
In the cases of either ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{s,e}\leq r_{1}}$ or ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{s,e}\geq r_{N}}$, it is apparent that any value of $d_{s,e}$ guarantees isolation, as the left and right expanding intervals will not include any other change-point, besides $r_{1}$ and $r_{N}$, respectively, before detection occurs. Considering now the more complicated case, where the location of the largest difference lies between two change-points, define $\tilde{\delta}_{j}$ as in (14) to be the distance between the consecutive change-points $r_{j}$ and $r_{j+1}$. In addition, $\delta_{s,e}^{j}=\min\{r_{j+1}-d_{s,e},d_{s,e}-r_{j}\}$ for $j=0,1,\ldots,N$. Without loss of generality, suppose that $d_{s,e}\in\left[r_{J},r_{J+1}\right]$ for some $J\in\{1,\ldots,N-1\}$. Requiring
$0\leq\delta_{s,e}^{J}\leq\frac{\tilde{\delta}_{J}}{2}-\frac{3\delta_{T}}{4n},$ (6)